OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 A. -7 B. 7 C. 9 D.-9

OpenStudy (ivettef365):

You multiply the first equation by 4 and the second equation by -5

OpenStudy (ivettef365):

Then you get: 20x +16y =4 -20x -15y = 5

OpenStudy (ivettef365):

When you subtract then you eliminate the x and you have y = 9

OpenStudy (anonymous):

thanks!!